-3=n over 6 + 9
be very detailed please

Answer by checkley71(8403)

-3=(n/6)+9
n/6=-3-9
n/6=-12 now cross multiply
n=-72 answer.
----------------------------------------------
or
-3=n/(6+9)
-3=n/15 cross multiply
n=-45 answer.